OpenStudy (anonymous):

Simplify. Express your answer in the form a + bi. (2+ i)(3 − 2i)−(i + 1)

OpenStudy (anonymous):

You can just distribute the first part. I'm going to use the classic FOIL method, just like you might do for factors of a quadratic or other polynomial. Let's look at this. (2+i)(3-2i) first: 2*3=6 outside: 2*-2i=-4i inside: i*3=3i last: i*-2i=-2i^2=2 6-4i+3i+2 8-i Now, we can subtract. (8-i)-(i+1) 8-i-i-1 7-2i

OpenStudy (anonymous):

i dont understand this step last: i*-2i=-2i^2=2

OpenStudy (anonymous):

2i^2 = 2 ? how ?